OpenStudy (anonymous):

if 10^x divided by 10^y =100, then x = ? a) y+2 b) y-2 c) 2-y d) 2y e) y/2

Parth (parthkohli):

\[10^{x}\div 10^y = 10^2\]\[\implies 10^{x - y} = 10^2\]\[\implies x - y = 2\]Hint provided.

Parth (parthkohli):

Solve for \(x\) in the above equation.

Parth (parthkohli):

@robtobey are you sure that is correct?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Sorry. x = y + 2

Parth (parthkohli):

Yes, that is correct. ^